The values of lattice energies (kJ/mol) for NaCl and AgCl are given below as calculated from
Born-Haber Cycle and Born-Lande equation, respectively.
U(B-H) U(B-L)
NaCl 787 751
AgCl 915 734
a) Why is the value for the lattice energy in each case, greater from Born-Haber cycle than that
from the Born-Lande equation?
b) What is the percentage difference between the two values for NaCl and AgCl? Rationalize
this on the basis of the difference in electronegativities of Na:Cl and Ag:Cl.
1
Expert's answer
2020-06-08T15:44:11-0400

Answer:

a) Born-Lande equation does not take into account the covalent component of the cation–anion bonds, as a result of which the obtained energies of the crystal lattices have underestimated values.

b) The percentage difference between the two values for NaCl and AgCl is:

for NaCl: "\frac{787-751}{787}=0.046=4.6%" %

for AgCl: "\frac{915-734}{915}=0.198=19,8" %

So, we can see that In the case of NaCl, the error of the approximate Born-Lange equation is small. This shows that NaCl is almost close to the ideal ionic compound. For AgCl, the error is large, since Ag atoms are much more electronegative then Na atoms. Accordingly, chemical bonds in the AgCl crystal lattice have a significant covalent component.